# Enabling Weak Client Participation via On-Device Knowledge Distillation in Heterogeneous Federated Learning

### v1

@@ 4 Experiments
CIFAR-10 32 (0.2) 30 400 100 $49.11\pm 0.1\%$ $59.21\pm 1.5\%$ $\textbf{62.43}\pm 0.1\%$
FEMNIST 20 (0.02) 30 200 100 $64.19\pm 0.1\%$ $53.75\pm 0.6\%$ $\textbf{67.44}\pm 0.1\%$
Figure 3: The validation accuracy curves corresponding to Table 3.
@@
We conduct performance evaluations with computer vision: CIFAR-10 (ResNet20 [31]), CIFAR-100 (ViT [32]), FEMNIST (CNN), natural language processing: IMDB (Bidirectional LSTM), and audio recognition: Google Speech Command (CNN) benchmarks. We report accuracy averaged across at least 3 independent runs. For all the methods in our comparative study, the algorithm-specific hyper-parameters were highly tuned using appropriate grid searches. See Appendix for the hyper-parameter settings in detail.
Non-IID Dataset Settings – We use label-based Dirichlet distributions ($\alpha=0.1$ which represents highly non-IID FL environments). FEMNIST is naturally non-IID, and we use 100 clients’ data as unlabeled local data. Google Speech is also naturally non-IID, but each local dataset is quite large. So, we use an artificial distribution to make it more strongly non-IID. We activate random $20\%$ of the total clients at every round. In all our experiments, a half of dataset is distributed to all clients as unlabeled local data.
@@
CIFAR-10 $57.98\pm 0.3\%$ $58.38\pm 0.1\%$ $58.84\pm 0.1\%$ $25.95\pm 0.1\%$ $61.13\pm 0.1\%$ $34.27\pm 0.1\%$ $\textbf{62.43}\pm 0.1\%$
FEMNIST $64.27\pm 0.1\%$ $64.80\pm 0.1\%$ $66.90\pm 0.2\%$ $56.71\pm 0.5\%$ $65.23\pm 0.1\%$ $65.06\pm 0.1\%$ $\textbf{67.44}\pm 0.1\%$
@@ 4.1 Performance Evaluation
We first define ’weak’ and ’strong’ clients based on their system capability. Here, weak client indicates the devices that have insufficient system resources to train the large target model. The strong client is the devices that have system resources strong enough to train the full target model. For extensive empirical study, we use a variety of different weak clients as shown in Table 4.
Table 3 shows the performance comparison between two baselines and the proposed method, and Figure 3 shows the corresponding validation accuracy curves (See Appendix for the loss curves). We consider two baselines: weak-only, where the model size is reduced to fit within all clients’ memory space and strong-only, where the weak clients are excluded and only the strong clients participate in the large model training. These two baselines are the most straightforward options for standard FL. By comparing our method to these two baselines, we can evaluate how well the proposed on-device KD utilizes the weak clients’ local data to strengthen the large target model. To ensure a fair comparison in terms of computational cost, the strong-only method performs $2\tau$ local steps, while our method runs $\tau$ local SGD steps and $\tau$ KD steps.
@@ 4.2 Comparative Study
We compare our method to three homogeneous FL methods: FedAvg, MOON, and FedDF, and heterogeneous FL methods: DS-FL, FedMD, and FedGEM [29] as shown in Table 5. The homogeneous FL methods use a small model that has $25\%$ of channels at each layer. The heterogeneous FL methods use the full model at $20\%$ of the clients and the small model at the rest of the clients. These settings well represent the realistic FL environments consisting of edge devices with heterogeneous system resources.
@@
Comparison to Homogeneous FL Methods – MOON achieves higher accuracy than FedAvg showing a practical use-case of contrastive loss. FedDF achieves even higher accuracy than MOON thanks to its server-side KD with public data. However, because these methods do not consider the system heterogeneity, they cannot train large models that do not fit into weak clients’ memory space. In contrast, our method can train large target models achieving superior performance by exploiting the strong clients’ system resources.
@@
Comparison to Heterogeneous FL Methods – DS-FL shows much lower accuracy than others due to the personalized local models. This result is well aligned with our ablation study shown in Figure 1 such that the logit ensemble harms the soft target quality dramatically losing the accuracy. FedMD achieves much higher accuracy than DS-FL. However, the performance gain mainly comes from the labeled public data shared across all the clients. In addition, FedMD is not applicable to the environments where the unlabeled data is distributed to the edge since the global logit should be calculated from the local logits obtained from the same data across all the clients. FedGEM achieves inconsistent performance across benchmarks. Due to the poor soft target quality caused by the logit ensemble, it severely loses the accuracy. FedGEM also assumes the labeled public data like FedMD, which is not realistic. Based on this comparative study, we can conclude that our method utilizes the unlabeled data more effectively than other methods.

### v2

@@ 4 Experiments
CIFAR-10 32 (0.2) 30 400 100 $57.47\pm 0.8\%$ $59.21\pm 1.5\%$ $\textbf{62.43}\pm 0.1\%$
FEMNIST 20 (0.02) 30 200 100 $64.27\pm 0.1\%$ $53.75\pm 0.6\%$ $\textbf{67.44}\pm 0.1\%$
Figure 3: The training loss (top) and validation accuracy (bottom) curves corresponding to Table 4.
@@
CIFAR-10 $57.47\pm 0.8\%$ $58.38\pm 0.1\%$ $41.62\pm 0.1\%$ $25.95\pm 0.1\%$ $61.13\pm 0.1\%$ $34.27\pm 0.1\%$ $\textbf{62.43}\pm 0.1\%$
FEMNIST $64.27\pm 0.1\%$ $4.80\pm 0.1\% 66$$6.90\pm 0.2\%$ $56.71\pm 0.5\%$ $65.23\pm 0.1\%$ $65.06\pm 0.1\%$ $\textbf{67.44}\pm 0.1\%$
@@
We conduct performance evaluations in computer vision: CIFAR-10 (ResNet20 [13]), CIFAR-100 (ViT [10]), FEMNIST (CNN), natural language processing: IMDB (Bidirectional LSTM), and audio recognition: Google Speech Command (CNN) benchmarks. We report accuracy averaged across at least 3 independent runs. For all the methods in our comparative study, the algorithm-specific hyper-parameters were highly tuned using appropriate grid searches. See Appendix for the hyper-parameter settings in detail
Non-IID Dataset Settings – We use label-based Dirichlet distributions ($\alpha=0.1$ which represents highly non-IID FL environments). FEMNIST is naturally non-IID, and we use 100 clients’ data as unlabeled local data. Google Speech is also naturally non-IID, but each local dataset is quite large. So, we use an artificial distribution to make it more strongly non-IID. We activate random $20\%$ of the total clients at every round. In all our experiments, half of the dataset is distributed to all clients as unlabeled local data.
@@ 4.1 Performance Evaluation
In all our experiments, we follow the definitions of strong and weak clients as discussed in Section 3.2. As shown in Table 3, we explore a variety of auxiliary model sizes for a comprehensive empirical study.
Table 4 shows the performance comparison between two baselines and the proposed method, and Figure 3 shows the corresponding training loss and validation accuracy curves, respectively. We consider two baselines: weak-only, where the model size is reduced to fit within all clients’ memory space and strong-only, where the weak clients are excluded and only the strong clients participate in the large model training. These two baselines are the most straightforward options for standard FL. By comparing our method to these two baselines, we can evaluate how well the proposed on-device KD utilizes the weak clients’ local data to strengthen the large target model. To ensure a fair comparison in terms of computational cost, the weak-only and strong-only methods perform $2\tau$ local steps, while our method runs $\tau$ local SGD steps and $\tau$ KD steps.
@@ 4.2 Comparative Study
We compare our method to three homogeneous FL methods: FedAvg, MOON, and FedDF, and heterogeneous FL methods: DS-FL, FedMD, and FedGEM [7], as shown in Table 5. The homogeneous FL methods use a small model that has $25\%$ of channels at each layer. The heterogeneous FL methods use the full model at $20\%$ of the clients and the small model at the rest of the clients. These settings well represent the realistic FL environments consisting of edge devices with heterogeneous system resources.
@@
Comparison to Homogeneous FL Methods – MOON achieves higher accuracy than FedAvg, demonstrating a practical use-case of contrastive loss. FedDF achieves even higher accuracy than MOON thanks to its server-side KD with public data. However, because these methods do not consider the system heterogeneity, they cannot train large models that do not fit into weak clients’ memory space. In contrast, our method can train large target models, achieving superior performance by exploiting the strong clients’ system resources.
@@
Comparison to Heterogeneous FL Methods – DS-FL shows much lower accuracy than others due to the personalized local models. This result is well aligned with our ablation study shown in Figure 1, indicating that the logit ensemble dramatically harms the soft target quality, leading to a loss in accuracy. FedMD achieves much higher accuracy than DS-FL. However, the performance gain mainly comes from the labeled public data shared across all the clients. In addition, FedMD is not applicable to environments where unlabeled data is distributed to the edge since the global logit should be calculated from the local logits obtained from the same data across all the clients. FedGEM achieves inconsistent performance across benchmarks. Due to the poor soft target quality caused by the logit ensemble, it severely loses the accuracy. FedGEM also assumes the labeled public data like FedMD, which is not realistic. Based on this comparative study, we can conclude that our method utilizes the unlabeled data more effectively than other methods.
4.3 Ablation Study
Auxiliary Model Size – To analyze the impact of the auxiliary model size on the target model’s performance, we conduct an ablation study as shown in Table 6. First, we set the number of strong clients to $20\%$ of the total clients. Then, we measure the accuracy of the auxiliary and target models using different auxiliary model sizes. One clear trend observed across all settings is that the target model accuracy improves as the auxiliary model size increases. This trend demonstrates that the auxiliary model more effectively transfers the global knowledge to the target model as its size grows.
One intriguing observation is that, for CIFAR-10 and Google Speech, the auxiliary model achieves even higher accuracy than the target model when its size is $75\%$. This is the case where the amount of data plays a more important role in achieving good accuracy than the model size difference. That is, the increased learning capability from a larger target model does not significantly improve performance, as the sufficiently strong auxiliary model can already learn global knowledge effectively. This result shows that the proposed method becomes more practical as the system becomes strongly heterogeneous.
Strong Client Ratio – We conduct another ablation study to analyze the impact of strong client ratio on FL performance. Table 7 shows the target model’s accuracy with a varying number of strong clients. The auxiliary model size is $25\%$ of the target model. The unlabeled data size is fixed regardless of the number of strong clients. In all the benchmarks, the target model performance is improved as the ratio of the strong clients increases. This result demonstrates the benefit of having an auxiliary model: As the ratio of the strong clients increases, the target model can learn from more unlabeled data, which results in improving its validation accuracy. Therefore, our method can be considered as a promising heterogeneous FL method which scales up more effectively.
Dataset Ratio Auxiliary Acc. Target Model Acc.
CIFAR-10 $25\%$ $49.11\pm 0.1\%$ $62.43\pm 0.1\%$
ResNet20 $50\%$ $60.86\pm 1.5\%$ $64.62\pm 0.5\%$
$75\%$ $6.65\pm 0.6\% 66$$6.39\pm 0.8\%$
FEMNIST $25\%$ $64.19\pm 0.1\%$ $67.44\pm 0.1\%$
CNN $50\%$ $65.89\pm 0.4\%$ $67.73\pm 0.2\%$
$75\%$ $66.67\pm 0.3\%$ $68.63\pm 0.3\%$
Google Speech $25\%$ $66.83\pm 0.3\%$ $71.95\pm 0.5\%$
CNN $50\%$ $74.54\pm 0.1\%$ $75.25\pm 0.1\%$
$75\%$ $76.77\pm 0.1\%$ $75.94\pm 0.2\%$
Table 6: The accuracy comparison across varying auxiliary model sizes. $20\%$ of the total clients are strong.
Dataset Strong / Total Clients Target Model Acc.
CIFAR-10 10 / 100 $64.37\pm 0.1\%$
ResNet20 20 / 100 $68.29\pm 0.7\%$
40 / 100 $70.20\pm 0.8\%$
FEMNIST 10 / 100 $66.36\pm 0.1\%$
CNN 20 / 100 $67.44\pm 0.1\%$
40 / 100 $67.78\pm 0.1\%$
Google Speech 10 / 100 $67.91\pm 0.1\%$
CNN 20 / 100 $71.95\pm 0.5\%$
40 / 100 $74.72\pm 0.4\%$
Table 7: The accuracy comparison across varying strong client ratios. The auxiliary model size is $25\%$.
@@ 5 Conclusion
Acknowledgments
This work was supported by INHA UNIVERSITY Research Grant.